What’s the answers form 33-40 answers fast put only the answers

for Sio,000 at 8% from one of its was unable to pay, and the 33) 33) On January 1, Southwest Company accepted a one-year note customers. When the note matured on December 31, the customer company Tecorded the dishonor. The amount of the debit recorded on December 31 D) $4800 A) 555,200 B) 560,000 C) 564,800 34) 34) Which of the following is included in the numerator of the acid-test ratio? Al canig cash equivalents, inventory, shon-term investments, net curment receivables B) total current assets C) cash inchuding cash equivalents, short-term investments, net current recrivables D) accounts receivable and inventory 35) 35) Logan Sales provides the following ๒formatiot Net credit sales: $790,000 Beginning net accounts receivable: $43,000 Ending net accounts receivable: $22,000 Calculate the accounts neceivable turmover ratio. (Round your answer to the nearest whole number.) D) 36 times B)34times C) 18 times A) 24 times 36 A) Plant assets ane long-lived tangible assets used in the operations of a business B) Deprediation is necoeded on all plant assets C) Plant assets not currently bering used in business operations are classified as long-term 36) Which of the following statements is incornect? D) The life cycle of a plant asset includes: acquisition, usage, and disposal 37) Which of the following should be included in the cost of land? A) oost of installing signage cot of intalling fences B) cost to bulld sidewalks on the land D) cost to clear the land of old buildings 8) Which of the following is NOT recorded as the cost of a building? B) contractor charges D) fending A) building permits andhitectural fees 39) Millbum Company has aquired a property that included both land and a building for $530,000. The company hired an appraiser who has determined that the market value of the land is $320,000 and that of the building is $480,000. At what amount should the company record the cost of land? Round any intermediate caloulations to two decimal places, and your final answer to the nearest dollar) A)5349.800 B) 5265,000 C) $212,000 D) $128000 A companys accountant capitalized a payment that should have been recorded as a nevenue expenditure How will this error affect the companys financial statements? A) Liabilities will be overstated
Answer #1

33)correct option is "C" - 64800

Amount of debit will be equal to amount of note receivable plus interest accrued for one year

            = 60000+ (60000*8%)

           = 60000+4800

            =64800

34)correct option is "C"

Acid test ratio is given by Total current asset (excluding inventory and prepaid expense) divided by current liabilities

so all other current asset will be included in numerator except inventory and prepaid expense

35)correct option is "A"

Accounts receivable turnover ratio = Net credit sales /average accounts receivables

            = 790000/ 32500

             = 24.3   [rounded 24]

**Average accounts receivable= [43000+22000]/2=32500

36)correct option is "C"

Even if plant asset is not used in business operations they are classified as tangible fixed asset and not as long term investment .